What was the initial speed of a car that required 13.7 m to stop in 4.00s?
1
Expert's answer
2012-09-25T11:05:20-0400
We know that

V = at,

where V is velocity, a is acceleration and t is the time of acceleration. Also we know that

L = at²/2 ==> a = 2L/t²,

where L is the stop distance. So, the initial speed of a car was

V = at = (2L/t²)*t = 2L/t = 2*13.7[m]/4.00[s] = 6.85 m/s.